OpenStudy (anonymous):

What is the slope between the two points (4, 14) and (2, 4)?

OpenStudy (anonymous):

(y2-y1)/(x2-x1)

OpenStudy (hoblos):

(14-4)/(4-2) = 10/2 = 5

OpenStudy (anonymous):

(4-14)/(2-4) = -10/-2 = 5

OpenStudy (anonymous):

5

OpenStudy (anonymous):

slope m = (14 -4) / (4 -2) = 10/ 2 = 5